Reconsider the fast-food restaurant visit discussed in chapter 1 (statistical quality control 10th Ed). What levels of quality would you consider acceptable for the family of four on each visit and annually? What probability of good quality on each meal component (assume 10 meal components in a typical restaurant meal) would be required in order to achieve these targets?
Answer #1

Solution:

p(good meal) = 10/100

=0.1

p(good meal for all four) = (0.1)^4

p(good meal for all four) =0.0001

Now p(bad experience) = 1- 0.0001

p(bad experience) = 0.9999